An Angry Bird™ was launched at the speed of 175 m/s into the air at an angle of 30° above the ground. As it reaches maximum height, a wind with air current speed of 300 m/s sent the ball perpendicularly off-course. What is the speed of the air current with respect to the Angry Bird™?


1
Expert's answer
2021-03-01T12:39:28-0500
"v_x=v\cos{30}"

The speed of the air current with respect to the Angry Bird:


"V=\sqrt{300^2+(175\cos{}30)^2}=336\frac{m}{s}"